OpenStudy (anonymous):

@lalayhow bout m+n=1000,0.05m+0.06n+57 @lalayhow bout m+n=1000,0.05m+0.06n+57 @Mathematics

OpenStudy (lalaly):

thats supposed to be an equal sign before the 57 right?

OpenStudy (anonymous):

yes sorry

OpenStudy (anonymous):

n= 700

OpenStudy (anonymous):

m =300

OpenStudy (lalaly):

ok m=1000-n substitute 100-n in m in the second equation 0.05(1000-n)+0.06n=57 50-0.05n+0.06n=57 0.01n=7 n=700

OpenStudy (lalaly):

m=1000-700 = 300

OpenStudy (anonymous):

god ur smart

OpenStudy (lalaly):

lol practice makes perfect ;)

OpenStudy (anonymous):

yea